# From small eigenvalues to large cuts, and Chowla’s cosine problem

### v1

@@ 13.1 Least eigenvalue version
edges. Furthermore, there are at most $n^{2-\varepsilon_{0}}$ edges of $G$ not in $G[X]$, so $G$ and $\widetilde{G}$ differ by at most $n^{2-\varepsilon_{0}}+n^{2-1/6}\leq n^{2-\varepsilon}$ edges. This finishes the proof. ∎
13.2 MaxCut version
Finally, we prove the MaxCut version of Theorem 13.5. Most of the proof of this theorem is identical to the proof of Theorem 13.5, but a substantial difficulty arises when one tries the adapt the proof of Lemma 13.4. Before, we start with a variant of Lemma 13.2.
Lemma 13.7.
Let $s\in(0,1/60)$ and $\delta>2s$, then there exists $\varepsilon>0$ such that the following holds. Let $G$ be a graph on $n$ vertices such that $\operatorname{surp}^{*}(G)\leq n^{1+s}$. Then there exists $X\subset V(G)$ such that the number of edges not in $G[X]$ is at most $n^{2-\varepsilon}$, and $G[X]$ can be partitioned into cliques of size $n^{1-\delta}$.
Proof.
The proof this is almost identical to the proof of Lemma 13.2. The only difference is that we cite Theorem 12.3 instead of Theorem 12.1. We omit further details. ∎
Recall that Lemma 13.4 states that a pair of cliques in a graph with large smallest eigenvalue must form a $\delta$-homogeneous pair for some very small $\delta$. The proof of this relied on finding a simple forbidden subgraph $H_{k}$. Unfortunately, for graphs of small surplus, we do not have such a simple forbidden structure anymore. Instead, we show that if $G$ contains two large cliques, then the bi-adjacency matrix of the bipartite graph between them is close to a rank-1 matrix. Then we show that this implies that the bipartite graph is close to a graph of rank one, which is just a complete bipartite graph. We prepare the proof of this with a couple of lemmas.
Lemma 13.8.
Proof.
$\|u\|_{2}\|v\|_{2}\leq\|A\|_{F}+\sqrt{\delta}n\leq 2n.$
We may rescale $u$ and $v$ such that $\|u\|_{2}=\|v\|_{2}\leq\sqrt{2n}$. Let $\alpha=\delta^{1/6}$, and define $x,y\in\{0,1\}^{n}$ such that
$x(i)=\begin{cases}1&\mbox{ if }u(i)\geq\alpha\\ 0&\mbox{ otherwise,}\end{cases}$
and similarly
$y(i)=\begin{cases}1&\mbox{ if }v(i)\geq\alpha\\ 0&\mbox{ otherwise.}\end{cases}$
We show that $xy^{T}$ is a good approximation of $A$. Note that $\|A-xy^{T}\|_{F}^{2}$ is the number of pairs $(i,j)$ such that $A_{i,j}\neq x_{i}y_{j}$. We count these pairs in three cases.
Case 1.
$A_{i,j}=1$ and $x_{i}=0$.
In this case, we have $u_{i}<\alpha$. If $v_{j}\leq 1/(2\alpha)$, then $(A_{i,j}-u_{i}v_{j})^{2}>1/4$, so there are at most $4\delta n^{2}$ such pairs $(i,j)$. On the other hand, the number of $j$ such that $v_{j}\geq 1/(2\alpha)$ is at most $4\alpha^{2}n$, as $\|v\|_{2}^{2}=\sum_{j=1}^{n}v_{j}^{2}\leq 2n$. Therefore, the number of $(i,j)$ such that $A_{i,j}=1$ and $x_{i}=0$ is at most
$4\delta n^{2}+4\alpha^{2}n^{2}=4\delta n^{2}+4\delta^{1/3}n^{2}=O(\delta^{1/3}n^{2}).$
Case 2.
$A_{i,j}=1$ and $y_{j}=0$.
Case 3.
$A_{i,j}=0$ and $x_{i}=y_{j}=1$.
In this case, $u_{i}\geq\alpha$ and $v_{j}\geq\alpha$, so $(A_{i,j}-u_{i}v_{j})^{2}\geq\alpha^{4}$. Thus, the total number of pairs $(i,j)$ in this case is at most $\delta n^{2}/\alpha^{4}=\delta^{1/3}n^{2}$.∎
Next, we prove a simple technical lemma which shows that the union of two cliques has large surplus as long as the two cliques are not too disjoint, and do not overlap too much.
Lemma 13.9.
Let $G$ be a graph such that $V(G)=C_{1}\cup C_{2}$ and $E(G)=\binom{C_{1}}{2}\cup\binom{C_{2}}{2}$. Let $|C_{1}\setminus C_{2}|=a$, $|C_{2}\setminus C_{1}|=b$ and $|C_{1}\cap C_{2}|=c$. Then
$\operatorname{surp}(G)\geq\frac{1}{4}\min\{a^{2},b^{2},c^{2}\}.$
Proof.
Let $A=C_{1}\setminus C_{2}$, $B=C_{2}\setminus C_{1}$, and $C=C_{1}\cap C_{2}$. We may assume that $a=b$. Otherwise, if, say $a\leq b$, we remove vertices of $B\setminus C$ until its size is exactly $a$. Then it is enough to show that the resulting graph has surplus at least $\frac{1}{4}\min\{a^{2},c^{2}\}$.
The number of edges of $G$ is
$2\binom{a+c}{2}-\binom{c}{2}<a^{2}+2ac+\frac{c^{2}}{2}.$
If $c\leq a$, then define the cut $(U,V)$ such that $U$ is some $(a+c)/2$ element subset of $A$ together with some $(a+c)/2$ element subset of $B$. The number of edges in this cut is $(a+c)^{2}/2$. Hence, the surplus of $G$ is at least $\frac{1}{2}(a+c)^{2}-\frac{1}{2}e(G)=c^{2}/4$.
If $c\geq a$, then define the cut $(U,V)$ such that $U$ is some $(a+c)/2$ element subset of $C$. Then the number of edges in this cut is $\frac{a+c}{2}\cdot\frac{c-a}{2}+2\frac{a+c}{2}\cdot a=\frac{c^{2}}{4}+\frac{3}{4}a^{2}+ac$. Therefore, the surplus is at least $a^{2}/4$. ∎
Now we are ready to prove our lemma about the surplus of the complement of bipartite graphs.
Lemma 13.10.
Let $s$ and $\delta>0$ such that $s+6\delta<1/2$. Let $H$ be a bipartite graph with vertex classes of size $n$, and let $G=\overline{H}$. If $\operatorname{surp}^{*}(G)\leq n^{1+s}$, then either $e(H)\leq n^{2-\delta}$ or $e(H)\geq n^{2}-n^{2-\delta}$.
Proof.
Let $A$ be the adjacency matrix of $G$ with eigenvalues $\lambda_{1}\geq\dots\geq\lambda_{2n}$. Furthermore, let $M$ be the adjacency matrix of $H$, and let $\mu_{1}\geq\dots\geq\mu_{2n}$ be the eigenvalues of $M$. As $H$ is bipartite, we have $\mu_{i}=-\mu_{2n+1-i}$ for $i\in[2n]$. By Lemma 5.3 (ii) and Lemma 11.1, we have
$\operatorname{surp}^{*}(G)\geq\Omega\left(\frac{1}{\sqrt{n}}\sum_{\lambda_{i}<0}\lambda_{i}^{2}\right)\geq\Omega\left(\frac{1}{\sqrt{n}}\sum_{i\neq 1,\mu_{i}>0}\mu_{i}^{2}\right)=\Omega\left(\frac{1}{\sqrt{n}}\sum_{i\neq 1,2n}\mu_{i}^{2}\right).$
Hence, if $\operatorname{surp}^{*}(G)\leq n^{1+s}$, then we have $\sum_{i\neq 1,2n}\mu_{i}^{2}=O(n^{3/2+s})$.
On the other hand, we can express $\sum_{i\neq 1,2n}\mu_{i}^{2}$ as follows. The matrix $M$ has the form $M=\begin{pmatrix}0&B\\ B^{T}&0\end{pmatrix}$ with an appropriate $n\times n$ matrix $B$. Let $v_{1}$ be the principal eigenvector of $M$, then we can write $v_{1}=(u,v)$, where $u,v\in\mathbb{R}^{n}$ correspond to the two vertex classes of $H$. Then the eigenvector corresponding to the smallest eigenvalue $\lambda_{2n}=-\lambda_{1}$ is $v_{2n}=(u,-v)$, and we have
$\sum_{i\neq 1,2n}\mu_{i}^{2}$ $=\|M-\lambda_{1}v_{1}v_{1}^{T}-\lambda_{2n}v_{2n}v_{2n}^{T}\|_{F}^{2}$
$=\left\|\begin{pmatrix}0&B\\ B^{T}&0\end{pmatrix}-\lambda_{1}\begin{pmatrix}uu^{T}&uv^{T}\\ vu^{T}&vv^{T}\end{pmatrix}+\lambda_{1}\begin{pmatrix}uu^{T}&-uv^{T}\\ -vu^{T}&vv^{T}\end{pmatrix}\right\|_{F}^{2}=2\|B-2\lambda_{1}uv^{T}\|_{F}^{2}.$
Therefore, $\|B-2\lambda_{1}vu^{T}\|_{F}=O(n^{3/2+s})$. But then by Lemma 13.8, there exist $x,y\in\{0,1\}^{n}$ such that $\|B-xy^{T}\|_{F}^{2}=O(n^{5/6+s/3})$. The matrix $xy^{T}$ corresponds to a complete bipartite graph between the vertex classes of $H$, let $\widetilde{H}$ denote this complete bipartite graph, and let $X_{0}$ and $Y_{0}$ denote its vertex classes. Note that $e(\widetilde{H})=\|xy^{T}\|_{F}^{2}=|X_{0}||Y_{0}|$, and $\|B-xy^{T}\|_{F}^{2}$ is the number of edges $\widetilde{H}$ differs from $H$. Therefore, if $e(\widetilde{H})\leq n^{2-\delta}/2$, then $e(H)\leq e(\widetilde{H})+\|B-xy^{T}\|_{F}\leq n^{2-\delta}$, so we are done. We can proceed similarly if $e(\widetilde{H})\geq n^{2}-n^{2-\delta}/2$. Hence, we may assume that $n^{2-\delta}/2\leq e(\widetilde{H})\leq n^{2}-n^{2-\delta}/2$. We show that this is impossible, by deriving that the surplus of $G$ is too large in this case.
Let $\widetilde{G}$ be the complement of $\widetilde{H}$. Then $\widetilde{G}$ and $G$ differ by at most $O(n^{5/6+s/3})$ edges. On the other hand, $\widetilde{G}$ is the union of two cliques, having vertex sets $C_{1}$ and $C_{2}$, where $X_{0}=C_{1}\setminus C_{2}$, $Y_{0}=C_{2}\setminus C_{1}$, and $C_{1}\cap C_{2}=V(G)\setminus(X_{0}\cup Y_{0})$. As $n^{2-\delta}/2\leq e(\widetilde{H})=|X_{0}||Y_{0}|$, we have $|X_{0}|,|Y_{0}|\geq n^{1-\delta}/2$. Also, as $e(\widetilde{H})\leq n-n^{2-\delta}/2$, we have $|C_{1}\cap C_{2}|=|V(G)\setminus(X_{0}\cup Y_{0})|\geq n^{2-\delta}/2$. Hence, by applying Lemma 13.9, we get that $\operatorname{surp}^{*}(\widetilde{G})\geq\Omega(n^{2-2\delta})$. But as $G$ and $\widetilde{G}$ differ by less than $O(n^{5/6+s/3})$ edges, and $5/6+s/3<2-2\delta$, this gives
$\operatorname{surp}^{*}(G)\geq\operatorname{surp}^{*}(\widetilde{G})-O(n^{5/6+s/3})\geq\Omega(n^{2-2\delta})>n^{1+s}$
as well, contradiction. ∎
Now we are ready to prove the main theorem of this section, that is, Theorem 1.4, which we restate here for the reader’s convenience. The proof is essentially the same as the proof of Theorem 13.5.
Theorem 13.11.
Let $s\in(0,1/60)$, then there exists $\varepsilon>0$ such that the following holds. Let $G$ be an $n$-vertex graph such that $\operatorname{surp}^{*}(G)\leq n^{1+s}$. Then $G$ is $n^{-\varepsilon}$-close to the disjoint union of cliques.
Proof.
The proof is almost identical to the proof of Theorem 13.5, but we use Lemmas 13.7 and 13.10 instead of Lemmas 13.2 and 13.4. We omit further details. ∎
Finally, we show that the previous theorem indeed implies Theorem 1.8 after complementation.
Proof of Theorem 1.8.
Define the positive discrepancy of a graph $G$ of edge density $p$ as
and define the negative discrepancy as
It was proved in [64], Lemma 2.6, that if $G$ is regular, then $\operatorname{surp}(G)=\Theta(\operatorname{disc}^{-}(G))$ and $\operatorname{bw}(G)=e(G)/2-\Theta(\operatorname{disc}^{+}(G))$. Moreover, $\operatorname{disc}^{+}(G)=\operatorname{disc}^{-}(\overline{G})$. Therefore, the theorem follows from Theorem 1.4 after taking complements, and noting that if a regular graph $G$ is close to a complement of a disjoint union of cliques, then $G$ is close to a Turán graph. ∎
14 Concluding remarks — Chowla’s cosine problem in finite groups
@@
Littlewood’s $L_{1}$-problem and Chowla’s cosine problem have also been studied in the setting of groups. Green and Konyagin [42] proposed to study the smallest $L_{1}$-norm of a dense set $A\subseteq\mathbb{Z}/p\mathbb{Z}$. If $\widehat{\mathds{1}_{A}}$ denotes the Fourier transform of the indicator function of $A$ over $\mathbb{Z}/p\mathbb{Z}$, they showed that $\sum_{r}|\widehat{\mathds{1}_{A}}(r)|\geq(\log p)^{1/3-o(1)}$, which was later improved by Sanders [70] to $(\log p)^{1/2-o(1)}$. For sparser sets $A\subseteq\mathbb{Z}/p\mathbb{Z}$, a similar question has been studied by Schoen [72] and Konyagin and Shkredov [55].
@@
Paralleling the extensions of the Littlewood $L_{1}$-problem, Sanders [71] extended Chowla’s problem to finite abelian groups $G$ as follows. For a symmetric subset $A\subset G$, one can define
$M_{G}(A)=\sup_{y\in G}-\widehat{\mathds{1}_{A}}(y),$
@@
where $\widehat{f}:G\rightarrow\mathbb{C}$ is the discrete Fourier transform of a function $f:G\rightarrow\mathbb{C}$, and $\mathds{1}_{A}$ is the indicator function of $A$. Since $A$ is a symmetric set, $\widehat{\mathds{1}_{A}}$ is a real function. The natural analogue of Chowla’s problem is to estimate the minimum of $M_{G}(A)$ over all centrally symmetric subsets $A$ of size $n$. One can quickly observe that $M_{G}(A)$ need not go to infinity with the size of $A$. Indeed, if $A$ is a subgroup of $G$, then $M_{G}(A)=0$. On the other hand, Sanders ([71, Theorem 1.3]) proved that if $A$ is far from a subgroup of $G$, then $M_{G}(A)$ is necessarily large. Formally, he proved that for every $\delta>0$ there exists $c(\delta)>0$ such that if $M_{G}(A)\leq|G|^{c(\gamma)}$, then there is some subgroup $H<G$ satisfying $|H\triangle A|\leq\delta|G|$. Noting that the image of the Fourier transform $\widehat{\mathds{1}_{A}}$ is the spectrum of the Cayley graph $\mbox{Cay}(G,A)$ generated by $A$, we can use our main results to immediately improve these bounds, and to extend them to non-abelian groups as well.
@@
If $G$ is a finite group and $A\subset G$ such that $A=A^{-1}$, define $M_{G}(A)=\max-\lambda$, where the maximum is taken among all eigenvalues of the Cayley graph $\mbox{Cay}(G,A)$, which then coincides with the earlier definition for finite abelian groups. We show that $M_{G}(A)$ is small if and only if $A$ is close to a subgroup of $G$.
Theorem 14.1.
@@
Let $\delta,\varepsilon>0$, then the following holds for every sufficiently large finite group $G$. Let $A\subset G$ such that $A=A^{-1}$. If $M_{G}(A)\leq|G|^{1/4-\varepsilon},$ then there exists a subgroup $H<G$ such that
$|H\triangle A|\leq\delta|G|.$
@@
Moreover, if $\alpha>0$ is sufficiently small with respect to $\varepsilon$ and $M_{G}(A)<|G|^{1/6-\varepsilon}$, then there exists a subgroup $H<G$ such that
$|H\triangle A|\leq|G|^{1-\alpha}.$
@@
The main idea of the proof is to show that $\mbox{Cay}(G,A)$ is close to the disjoint union of cliques if and only if $A$ is close to a subgroup of $G$. This is proved in the following lemma.
Lemma 14.2.
@@
Let $G$ be a group and let $\subset G, A=A$$A^{-1}$, such that the number of pairs $(x,y)\in A\times A$ such that $xy^{-1}\not\in A$ is at most $\varepsilon|A|^{2}$. Then there exists a subgroup $H<G$ such that $|H\triangle A|\leq O(\varepsilon^{1/2}|A|)$.
@@
In the proof, we use an old theorem of Freiman [37] on sets of very small doubling, sometimes referred to as Freiman’s $3/2$-theorem. See also the blog of Tao [73] for a short proof. Given subsets $A,B$ of a group $G$, we write $AB:=\{xy:x\in A,y\in B\}$.
Lemma 14.3 (Freiman’s $3/2$-theorem).
@@
Let $G$ be a group and let $A\subset G$ such that $|AA^{-1}|\leq\frac{3}{2}|A|$. Then $AA^{-1}$ and $A^{-1}A$ are both subgroups of $G$.
Proof of Lemma 14.2.
@@
We may assume that $\varepsilon<1/1000$, otherwise the statement is trivial. Also, we can assume that the identity $1_{G}\in A$, as adding $1_{G}$ does not change the number of pairs $(x,y)\in A\times A$ with $xy^{-1}\not\in A$, and it only changes the size of $A$ by $1$.
Let $N$ be the number of pairs $(x,y)\in A\times A$ such that $xy^{-1}\not\in A$. Also, for every $x\in A$, let
$N(x)=|(xA)\triangle A|.$
Then using $A=A^{-1}$, we have
$N=\frac{1}{2}\sum_{x\in A}N(x).$
@@
Then by simple averaging, we have $|B|\geq(1-2\varepsilon/\delta)|A|=(1-\delta)|A|$. We also note that $B=B^{-1}$ as $N(x)=N(x^{-1})$, and $1_{G}\in B$. Observe that for every $x_{1},x_{2}\in B$, we can use the triangle inequality to write
@@
Therefore,
$(1-\delta)|A|\leq|B|\leq|B\cdot B|\leq(1+4\delta)|A|,$
which also implies that $|B\cdot B|\leq 3|B|/2$. Hence, we can apply Lemma 14.3 to conclude that $B\cdot B$ is a subgroup of $G$. Furthermore, as $1_{G}\in B$, we have $B\subset B\cdot B$, so $|A\cap B\cdot B|\geq|B|$. In conclusion
@@
With Lemma 14.2 in our hands, Theorem 14.1 follows almost immediately from Theorems 9.4 and 13.5.
Proof of Theorem 14.1.
@@
We start with the first part of the theorem. We may assume that $|A|\geq\delta|G|$, otherwise the statement is trivial by choosing $H=\{1_{G}\}$. Let $\Gamma=\mbox{Cay}(G,A)$, and let $\lambda_{n}=-M_{G}(A)$ be the smallest eigenvalue of $\Gamma$, $n=|G|$. We may assume that $1_{G}\not\in A$, by noting that removing $1_{G}$ shifts the eigenvalues by -1. Therefore, $\Gamma$ is a simple graph with no self loop. By Theorem 9.4, the inequality $|\lambda_{n}|\leq n^{1/4-\varepsilon}$ implies that $\Gamma$ is $\beta$-close to a disjoint union of cliques for any $\beta>0$, given $n>n_{0}(\beta,\varepsilon)$. But then $\Gamma$ contains at most $3\beta n^{3}$ cherries by Lemma 9.3. This implies that there are at most $6\beta n^{2}\leq 6(\beta/\delta^{2})|A|^{2}$ pairs $(x,y)\in A\times A$ such that $xy^{-1}\not\in A$. By Lemma 14.2, then $|A\triangle H|\leq O(\beta^{1/2}/\delta|A|)\leq\delta n$ for some subgroup $H<G$, assuming $\beta\ll\delta^{4}$.

### v2

@@ 11 Edit distance from the union of cliques
edges. Furthermore, there are at most $n^{2-\alpha_{0}}$ edges of $G$ not in $G[X]$, so $G$ and $\widetilde{G}$ differ by at most $n^{2-\alpha_{0}}+n^{11/6}\leq n^{2-\alpha}$ edges. This finishes the proof. ∎
12 Further directions
We conclude our paper by discussing some open problems. One of our main contributions is Theorem˜1.2, which shows that every graph of average degree $d$ with $|\lambda_{n}|\leq d^{\gamma}$, where $\gamma\in(0,1/10)$, contains a clique of size $d^{1-4\gamma}$. The constants $1/10$ and $4$ are likely not optimal, and it would be interesting to understand the limits of this theorem. Yang and Koolen [86] conjecture that if $d$ is exponentially larger than $\lambda_{n}$, then $G$ contains a clique of size $\Omega(d/|\lambda_{n}|^{2})$. We believe this can be already true if $|\lambda_{n}|=O(d^{\gamma})$ for some small $\gamma$.
Question 12.1.
Let $G$ be an $n$-vertex graph of average degree $d$ and $|\lambda_{n}|=\Theta(d^{\gamma})$ for some $\gamma\in(0,1/2)$. How small can its clique number $\omega(G)$ be?
Moreover, despite the substantial progress presented in this paper, finding precise exponents for the MaxCut problem in $H$-free graphs remains open. In particular, it is still an intriguing problem to determine the largest exponent $\alpha_{r}$ such that every $K_{r}$-free graph $G$ with $m$ edges satisfies $\operatorname{surp}(G)\geq\Omega(m^{\alpha_{r}})$. The work of Alon [2] shows that $\alpha_{3}=4/5$, but the value of $\alpha_{4}$ is already unknown. Following our calculations, we established that $\alpha_{r}>0.51$ for all fixed $r\geq 3$, improving the barrier $1/2$. However, the celebrated conjecture of Alon, Bollobás, Krivelevich and Sudakov [3] asserts that $\alpha_{r}>3/4$.
Acknowledgements
We are grateful to Ilya Shkredov for bringing Chowla’s problem to our attention, and pointing out connections to our earlier results. Moreover, we thank Benjamin Bedert for sharing his manuscript on Chowla’s problem and for an insightful conversation about it. Finally, we would like to thank Igor Balla, Clive Elphick, Victor Falgas-Ravry, Jacob Fox, Lianna Hambardzumyan, Jack Koolen, Anqi Li, Nitya Mani, Benny Sudakov, and Quanyu Tang for many valuable discussions on various parts of this project.
Appendix A Chowla’s cosine problem in finite groups
@@
In Section 1.1 we discussed Chowla’s cosine problem and its relation to Littlewood’s $L_{1}$-problem over the integers. Variants of these two problems have also been studied in the setting of discrete abelian groups. Namely, in 2009, Green and Konyagin [49] studies how small the $L_{1}$-norm of a dense set $A\subseteq\mathbb{Z}/p\mathbb{Z}$ can be. If $\widehat{\mathds{1}_{A}}$ denotes the Fourier transform of the indicator function of $A$ over $\mathbb{Z}/p\mathbb{Z}$, they showed that $\sum_{r}|\widehat{\mathds{1}_{A}}(r)|\geq(\log p)^{1/3-o(1)}$, which was later improved by Sanders [80] to $(\log p)^{1/2-o(1)}$. For sparser sets $A\subseteq\mathbb{Z}/p\mathbb{Z}$, a similar question has been studied by Schoen [82] and Konyagin and Shkredov [63].
@@
Paralleling the extensions of the Littlewood $L_{1}$-problem, Sanders [81] extended Chowla’s problem to discrete abelian groups $\Gamma$ as follows. For a symmetric subset $A\subset\Gamma$, one can define
$M_{\Gamma}(A)=\sup_{y\in\widehat{\Gamma}}-\widehat{\mathds{1}_{A}}(y),$
where $\widehat{\Gamma}$ is the dual group of $\Gamma$. Note that since $A$ is a symmetric set, $\widehat{\mathds{1}_{A}}$ is a real function.
@@
In this language, Chowla’s cosine problem asks to show that for every $A\subseteq\mathbb{Z}$ we have $M_{\mathbb{Z}}(A)\geq\Omega(\sqrt{|A|})$, but it is also natural to ask how small $M_{\Gamma}(A)$ can be in a general group $\Gamma$. One can quickly observe that $M_{\Gamma}(A)$ need not go to infinity with the size of $A$. Indeed, if $A$ is a subgroup of $\Gamma$, then $M_{\Gamma}(A)=0$. On the other hand, Sanders ([81, Theorem 1.3]) proved that if $A$ is far from a subgroup of $\Gamma$, then $M_{\Gamma}(A)$ is necessarily large. Formally, he proved that for every $\delta>0$ there exists $c(\delta)>0$ such that if $M_{\Gamma}(A)\leq|\Gamma|^{c(\delta)}$, then there is some subgroup $H<\Gamma$ satisfying $|H\triangle A|\leq\delta|\Gamma|$. Noting that the Fourier coefficients of $\mathds{1}_{A}$ correspond to the eigenvalues of the Cayley graph $\operatorname{Cay}(\Gamma,A)$ generated by $A$, we can use our main results to improve these bounds when $\Gamma$ is finite, and to extend them to non-abelian groups as well.
@@
Recall, the Cayley graph $\operatorname{Cay}(\Gamma,A)$, where $\Gamma$ is a finite group and $A\subseteq\Gamma$ is symmetric, has vertex set $\Gamma$ and two vertices $x,y\in\Gamma$ are adjacent if and only if $xy^{-1}\in A$. We define $M_{\Gamma}(A):=\max-\lambda$, where the maximum is taken among all eigenvalues of $\operatorname{Cay}(\Gamma,A)$, which then coincides with the earlier definition for finite abelian groups. As discussed above, $M_{\Gamma}(A)=0$ if $A$ is a subgroup of $\Gamma$, so $M_{\Gamma}(A)$ is small if $A$ is close to a subgroup of $\Gamma$. Here, we show the opposite: if $M_{\Gamma}(A)$ is small, then $A$ must be close to a subgroup of $\Gamma$.
Theorem A.1.
@@
Let $\delta,\gamma>0$, then the following holds for every sufficiently large finite group $\Gamma$. Let $A\subset\Gamma$ such that $A=A^{-1}$. If $M_{\Gamma}(A)\leq|\Gamma|^{\gamma}$ and $\gamma\in(0,1/4)$ then there exists a subgroup $H<\Gamma$ such that
$|H\triangle A|\leq\delta|\Gamma|.$
@@
Moreover, if $\alpha>0$ is sufficiently small as a function of $\gamma$ and $M_{\Gamma}(A)<|\Gamma|^{\gamma}$ for $\gamma\in(0,1/6)$, then there exists a subgroup $H<\Gamma$ such that
$|H\triangle A|\leq|\Gamma|^{1-\alpha}.$
@@
The main idea of the proof is to show that $\operatorname{Cay}(\Gamma,A)$ is close to the disjoint union of cliques if and only if $A$ is close to a subgroup of $\Gamma$. This is proved in the following lemma.
Lemma A.2.
@@
Let $\Gamma$ be a group and let $\subset\Gamma, A=A$$A^{-1}$, such that the number of pairs $(x,y)\in A\times A$ such that $xy\not\in A$ is at most $\varepsilon|A|^{2}$. Then there exists a subgroup $H<\Gamma$ such that $|H\triangle A|\leq O(\varepsilon^{1/2}|A|)$.
@@
In the proof, we use an old theorem of Freiman [43] on sets of very small doubling, sometimes referred to as Freiman’s $3/2$-theorem. See also the blog of Tao [83] for a short proof. Given subsets $A,B$ of a group $\Gamma$, we write $A\cdot B=AB=\{xy:x\in A,y\in B\}$.
Theorem A.3 (Freiman’s $3/2$-theorem).
@@
Let $\Gamma$ be a group and let $A\subset\Gamma$ such that $|AA^{-1}|<\frac{3}{2}|A|$. Then $AA^{-1}$ and $A^{-1}A$ are both subgroups of $\Gamma$.
Proof of Lemma A.2.
@@
We may assume that $1/|A|\leq\varepsilon<1/1000$, otherwise the statement is trivial. Also, we can assume that the identity $1_{\Gamma}\in A$, as adding $1_{\Gamma}$ does not increase the number of pairs $(x,y)\in A\times A$ with $xy^{-1}\not\in A$, and it only changes the size of $A$ by $1$.
Let $N$ be the number of pairs $(x,y)\in A\times A$ such that $xy\not\in A$. Also, for every $x\in A$, let
$N(x)=|(xA)\triangle A|=|(xA)\backslash A|+|A\backslash(xA)|=|(xA)\backslash A|+|(x^{-1}A)\backslash A|.$
Hence, we have
$N=\sum_{x\in A}|(xA)\setminus A|=\frac{1}{2}\sum_{x\in A}N(x).$
@@
Then by simple averaging, we have $|B|\geq(1-2\varepsilon/\delta)|A|=(1-\delta)|A|$. We also note that $B=B^{-1}$ as $N(x)=N(x^{-1})$, and $1_{\Gamma}\in B$. Our goal is to show to apply Freiman’s 3/2-theorem to the set $B$, and so we now show that $|BB^{-1}|=|B\cdot B|\leq\frac{3}{2}|B|$.
Observe that for every $x_{1},x_{2}\in B$, we can use the triangle inequality to write
@@
Since $|B|\geq(1-\delta)|A|$ and $\delta$ is sufficently small, we conclude that $|B\cdot B|<3|B|/2$. By Theorem˜A.3, $B\cdot B^{-1}=B\cdot B$ is a subgroup of $\Gamma$. Since $1_{\Gamma}\in B$, we have $B\subset B\cdot B$, so $|A\cap(B\cdot B)|\geq|B|$. In conclusion
@@
With Lemma˜A.2 in our hands, Theorem˜A.1 follows almost immediately from Theorems˜1.4 and 1.5.
Proof of Theorem˜A.1.
We start with the first part of the theorem. We may assume that $|A|\geq\delta|\Gamma|$, otherwise the statement is trivial by choosing $H=\{1_{\Gamma}\}$. Also, fix a parameter $\beta\ll\delta^{4}$.
@@
Let $G=\operatorname{Cay}(\Gamma,A)$, and let $\lambda_{n}=-M_{\Gamma}(A)$ be the smallest eigenvalue of $G$, $n=|\Gamma|$. We may assume that $1_{\Gamma}\not\in A$, by noting that removing $1_{\Gamma}$ shifts the eigenvalues by 1-1. Therefore, $G$ is a simple graph with no loops. If $n$ is sufficiently large as a function of $\delta$, Theorem˜1.4 shows that the inequality $|\lambda_{n}|\leq n^{\gamma}$ for $\gamma\in(0,1/4)$ implies that $G$ is $\beta$-close to a disjoint union of cliques. But then $G$ contains at most $3\beta n^{3}$ induced cherries by Lemma˜8.4. Recall, a cherry is a triple of vertices $u,v,w\in G$ such that $uv,vw\in E(G)$, but $uw\notin E(G)$. Since $G=\operatorname{Cay}(\Gamma,A)$, every cherry corresponds to a pair $x=uv^{-1}\in A,y=vw^{-1}\in A$ for which we have $xy=uw^{-1}\notin A$, and each such pair $(x,y)\in A$ corresponds to exactly $n$ cherries. Hence, there are at most $6\beta n^{2}\leq 6(\beta/\delta^{2})|A|^{2}$ pairs $(x,y)\in A\times A$ such that $xy\not\in A$. By Lemma˜A.2, then $|A\triangle H|\leq O(\beta^{1/2}/\delta|A|)\leq\delta n$ for some subgroup $H<\Gamma$, since $\beta\ll\delta^{4}$.